What should be the price of a bond which does not make coupons payments, has a one thousand dollar face value, and nine-year maturity, when the required rate of return is 3.5%?

$722.65

$687.44

$703.19

$733.73

$781.20
Answer #1

Price of bond = Redemption value x PVIF(r%,n)

Redmption vale = face value = 1000$

PVIF(r%,n) = 1/(1+r)^n

here rate of return = 3.5%

n = no of years till maturity = 9

PVIF(3.5%,9) = 1/(1+3.5%)^9

=1/(1+0.035)^9

=1/(1.035)^9

=1/1.362897

=0.73373

Thus Price of bond = 1000 x 0.73373

=$733.73